After carefully reading all the available information, prepare a two page (double-spaced) essay and answer the following questions:
Assume that we have the following data:
C=100+0.50Y
Ip=100-20r
Mt=0.10Y
Ms=100-10r
M=80

a. Build the IS-LM function.
b. If we assume an increase in Investments by 100 units, calculate again the IS-LM functions.
c. The intersection of IS-LM functions defines four areas. Analyze the behavior of the markets for goods and money for each area.
1
Expert's answer
2016-03-19T05:52:04-0400
C=100+0.50Y, Ip=100-20r, Mt=0.10Y, Ms=100-10r, M=80.
a. The IS–LM functions show the relationship between interest rates and real output, in the goods and services market and the money market.
The IS curve is defined by the equation Y = C + I + G + NX, so Y = 100 + 0.5Y + 100 - 20r - 80, Y = 240 - 40r. The LM curve is defined by the equation Mt = Ms, so 100 - 10r = 0.1Y, Y = 1000 - 100r.
b. If we assume an increase in Investments by 100 units, then the IS function will shift to the right.
c. The intersection of IS-LM functions defines four areas, and the behavior of the markets for goods and money for each area will change.